(4)/(x+1)-(2)/(x-1)=(3x+6)/(x^(2)-1)
Answer by Cromlix(4381)

4/(x + 1) - 2/(x - 1) = (3x + 6)/(x^2 - 1)
Common denominator = x^2 - 1
This because x^2 - 1 expands to give
(x + 1)(x - 1)
4(x - 1) - 2(x + 1) = (3x + 6)
4x - 4 - 2x - 2 = 3x + 6
4x - 2x - 3x = 6 + 4 + 2
- x = 12
x = - 12
Hope this helps.
:-)